# Probing $\alpha$ clustering in $^{12}\mathrm{C}$ at CSR energies using the Jet AA Microscopic Transport Model

### v1

@@ IV Conclusion
We find that $\alpha$ clustering leads to a more compact participant configuration than the Woods–Saxon case, while global transverse size and eccentricity fluctuations exhibit only a weak sensitivity to clustering. The root-mean-square radii of the two configurations differ by less than a few percent, indicating that the observed effects are not driven by trivial changes in nuclear size, but rather by differences in the microscopic spatial organization of nucleons. At this beam energy, radial observables remain directly sensitive to such geometric compactness. In particular, the proton mean transverse momentum is enhanced for $\alpha$-clustered configurations, whereas pion observables show little response. The $\langle p_{T}\rangle$ fluctuations of protons tend to be smaller in the $\alpha$-clustered configuration than in the Woods–Saxon case, consistent with reduced transverse-radius fluctuations of participant nucleons in the clustered configuration.
The collective anisotropic response of the system is examined through the flow coefficients. We find that $\alpha$ clustering is associated with an enhancement of the mean flow magnitudes, $\langle v_{n}\rangle=\sqrt{\langle v_{n}^{2}\rangle}$, at large $N_{\mathrm{part}}$, consistent with modifications of the average initial geometry and its collective response. In contrast, the event-by-event fluctuation strength of individual flow harmonics remains small and is not clearly distinguishable within the present statistical precision. While symmetric cumulants constructed from the initial-state eccentricities exhibit a noticeable sensitivity to clustering, the corresponding correlations among final-state flow harmonics do not show a comparably clear separation between clustered and unclustered configurations, suggesting that information on initial-state fluctuations is only partially retained during the dynamical evolution.

### v2

@@ IV Conclusion
We find that $\alpha$ clustering leads to a more compact participant configuration than the Woods–Saxon case, while global transverse size and eccentricity fluctuations exhibit only a weak sensitivity to clustering. The root-mean-square radii of the two configurations differ by less than a few percent, indicating that the observed effects are not driven by trivial changes in nuclear size, but rather by differences in the microscopic spatial organization of nucleons. At this beam energy, radial observables remain directly sensitive to such geometric compactness. In particular, the proton mean transverse momentum is enhanced for $\alpha$-clustered configurations, whereas pion observables show little response. The ensemble-averaged $\langle p_{T}\rangle$ fluctuations of protons tend to be smaller in the $\alpha$-clustered configuration than in the Woods–Saxon case, consistent with reduced transverse-radius fluctuations of participant nucleons in the clustered configuration.
The collective anisotropic response of the system is examined through the flow coefficients. We find that $\alpha$ clustering is associated with an enhancement of the root-mean-square flow magnitudes, $v_{n}\{2\}=\sqrt{\langle v_{n}^{2}\rangle}$, at large $N_{\mathrm{part}}$, consistent with modifications of the average initial geometry and its collective response. In contrast, the ensemble-averaged fluctuation strength of individual flow harmonics remains small and is not clearly distinguishable within the present statistical precision. While symmetric cumulants constructed from the initial-state eccentricities exhibit a noticeable sensitivity to clustering, the corresponding ensemble-averaged correlations among final-state flow harmonics do not show a comparably clear separation between clustered and unclustered configurations, suggesting that information on initial-state fluctuations is only partially retained during the dynamical evolution.